Solve the following system of equations.
2x + 4y – 3z = –7
3x + y + 4z = –12
x + 3y + 4z = 4
please show work

Answer by ewatrrr(24785)

 
Hi,
Solve the following system of equations.
2x + 4y – 3z = –7
3x + y + 4z = –12
x + 3y + 4z = 4
D = -72, Dx = 432, Dy = -144, Dz = -72
x = -6, y = 2, z = 1